A university bookstore recently sold a wirebound graph-paper notebook for $2.50, and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $118.60. How many of each type were sold?
thank you.

Answer by checkley77(12844)

2.50X+2.30(50-X)=118.60
2.5X+115-2.30X=118.60
.20X=118.60-115
.20X=3.6
X=3.6/.2
X=18 IS THE NUMBER OF WIREBOUND NOTEBOOKS WEE SOLD.
50-18=32 IS THE NUMBER OF COLLEGE-RULED NOTEBOOKS WERE SOLD.
PROOF:
2.50*18+2.3*32=118.60
45+73.60=118.60
118.6=118.60